A car is traveling at 45 m/s. The driver slams on the brakes, and comes to a stop in 6.3 seconds. Find the acceleration of the car.
1
Expert's answer
2020-11-23T10:26:47-0500

We can find the acceleration of the car from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{0\ \dfrac{m}{s}-45\ \dfrac{m}{s}}{6.3\ s}=-7.14\ \dfrac{m}{s^2}."

The sign minus means that the car deccelerates.

Answer:

"a=-7.14\ \dfrac{m}{s^2}."